OpenStudy (anonymous):

The rational function f(x) = p(x) q(x) , where p and q are quadratic functions, has the following proper- ties: (a) zeros at 3 and -1; (b) vertical asymptotes x = 1 and x = 4; (c) a horizontal asymptote y = 4. Determine f(2).

OpenStudy (anonymous):

p(x)/q(x)

OpenStudy (blockcolder):

Let's incorporate all the clues. If f has 0's at 3 and -1, this means that x-3 and x+1 are factors of p(x). Therefore, p(x)=A(x+1)(x-3). If f has VA's at 1 and 4, this means that x-1 and x-4 are factors of q(x). Therefore, q(x)=B(x-1)(x-4). An HA at y=4 means that the ratio A/B is 4. From all of these clues, we gather that \[f(x)=\frac{4(x-3)(x+1)}{(x-1)(x-4)}\] and thus \[f(2)=\frac{4(-1)(3)}{(1)(-2)}=6\]